# QGFace: Quality-Guided Joint Training For Mixed-Quality Face Recognition

### v1

@@ Abstract
The quality of a face crop in an image is decided by many factors such as camera resolution, distance, and illumination condition. This makes the discrimination of face images with different qualities a challenging problem in realistic applications. However, most existing approaches are designed specifically for high-quality (HQ) or low-quality (LQ) images, and the performances would degrade for the mixed-quality images. Besides, many methods ask for pre-trained feature extractors or other auxiliary structures to support the training and the evaluation. In this paper, we point out that the key to better understand both the HQ and the LQ images simultaneously is to apply different learning methods according to their qualities. We propose a novel quality-guided joint training approach for mixed-quality face recognition, which could simultaneously learn the images of different qualities with a single encoder. Based on quality partition, classification-based method is employed for HQ data learning. Meanwhile, for the LQ images which lack identity information, we learn them with self-supervised image-image contrastive learning. To effectively catch up the model update and improve the discriminability of contrastive learning in our joint training scenario, we further propose a proxy-updated real-time queue to compose the contrastive pairs with features from the genuine encoder. Experiments on the low-quality datasets SCface and Tinyface, the mixed-quality dataset IJB-B, and five high-quality datasets demonstrate the effectiveness of our proposed approach in recognizing face images of different qualities.
Index Terms:
Face recognition, discriminative representations, contrastive regularization, mixed-quality face recognition.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Existing face recognition (FR) methods have achieved near-perfect performances on high-quality (HQ) face images [1, 2, 3]. However, their performances significantly degrade when being applied to low-quality (LQ) images. Some methods aim to improve performance in specific image domains, such as light field images [4] and near-infrared images [5] and hard to be applied to general LQ images. Some other works focus on the face recognition where LQ images are adopted for the matching between the given probe images and the gallery images [6, 7, 8, 9, 10, 11]. However, due to the domain gap between the HQ and the LQ images, the discriminability for HQ images is somewhat reduced. It is hard to effectively learn from both the LQ and the HQ images simultaneously. In reality, the quality of a face image used for recognition is influenced by numerous factors such as camera resolution, distance, occlusion, and illumination condition. For instance, a high-resolution smart phone camera could produce low-quality face crops when there are too many faces in the picture or the photographer focuses on the scenery in the background. Meanwhile, in the surveillance situations, the image qualities also vary when the distance is changing (the setting in SCface [12] dataset). Domain gaps also exist among these low-quality images. In some limited situations where the image qualities are fixed, it is possible to select the proper model according to the image qualities. However, for the applications such as photo albums and complicated surveillance systems with various cameras, the qualities of face images vary a lot. Existing methods only focusing on either HQ or LQ images are not suitable for such scenarios.
@@
The problems with the mixed-quality face recognition using a single feature extractor lie in the following aspects. First, due to the domain gap between the HQ and the LQ data, it is hard to learn from and inference with images of different qualities simultaneously. Most existing methods emphasize either the HQ or the LQ images, while miss the other one. Some quality-invariant methods [9, 11, 10, 13, 14] attempt to understand the LQ images with the guidance of a pretrained HQ model. This explicitly teaches the model to lean to the LQ images and thus the performance degrades for the HQ images. Besides, as the HQ images are encoded by a fixed model, the compromise between images of difference qualities is also under-optimized. Second, we argue that there is a semantic problem when applying classification-based method on the LQ images. The identities of faces are hard to predict due to the natural variations of their poses, ages, and makeups for the HQ images. Compared with the HQ images, the LQ images typically have fewer details, which makes it challenging for the model to accurately classify them. This confusion arises because the model may struggle to differentiate between natural variations in the images and disturbances caused by the lower image qualities. As a result, the LQ images usually stumble the performance of the feature extractor trained with the classification loss [15].
@@
To address the above problems, we propose a novel approach namely quality-guided joint training for mixed-quality face recognition (QGFace). Our goal is to learn both the HQ and the LQ face images simultaneously, and recognize face images without distinction of qualities with a single feature extractor. To effectively utilize the data of different qualities, we design a quality partitioning strategy and apply different supervision signals according to the qualities of the images. Specifically, for the HQ images, the classification-based method is applied as they have achieved near-perfect performance for face recognition [2, 15, 1, 16]. For the LQ images which could mislead the classification loss, we adopt self-supervised contrastive learning which learns the face images at the instance level and alleviates the requirements for the image details. By purposefully taking advantages of the classification methods and the contrastive learning, we get an encoder which is applicable for images with different qualities. Furthermore, with the joint training strategy, we only need a single feature extractor during training and evaluation. This makes our method highly extensible, which could serve as the base component for future works.
@@
Nevertheless, taking contrastive learning from the unsupervised learning into the joint training scenario is non-trivial. Face identification asks for searching the related identity of a given image. Classification-based methods contain thousands of image-identity pairs to learn discriminative features which could meet the above requirement. For contrastive learning, it usually takes a large number of sample pairs to improve the stability and the accuracy. Momentum queue is a common choice [17]. It generates representations by an extra encoder which is momentum-updated with the training encoder. However, it loses track of the model which is also guided by the classification supervision signal, and the produced features are whitened as the encoder keeps average of all the past model statuses. Inspired by BroadFace [18], we propose a novel proxy-updated real-time contrastive queue to supply positive pairs from the genuine model with a large number of negative samples which is dynamically updated. The proxy is represented by the corresponding weights of an identity in the classifier.
@@
  • We propose a quality-guided joint training strategy for face recognition (QGFace) on the mixed-quality images. To better understand the LQ images which could mislead the classification methods, we partition the images into the HQ and the LQ subsets. Classification loss and contrastive loss are then applied on the two subsets respectively. In this way, we can comprehensively learn from both the LQ and the HQ faces to train a single encoder.
  • To support the feature search in face identification, we propose a proxy-updated real-time contrastive queue to ensure the discriminability of the LQ images learned by contrastive learning. The identity proxy representations are taken to update the queue features encoded in the previous training step.
  • We conduct extensive experiments on high-quality, low-quality, and mixed-quality datasets. Our approach achieves competitive results on different categories of datasets compared with the SOTA methods, which demonstrates the effectiveness of our QGFace.
Figure 1: Conventional pure classification method vs. our QGFace. (a) An FR training pipeline with a single classification loss. All the input images are taken into a single main loss. (b) Our pipeline additionally takes contrastive loss for low-quality images. We apply data augmentation on images and split their features into two parts: HQ and LQ ones. The HQ images are supervised by the classification loss, while the LQ images and their related HQ images are sent into the contrastive function. The gradient flow of HQ data is stopped in contrastive learning. A real-time queue is designed to provide an effective feature queue and support the large-scale feature comparison.
@@ II-A Encoder-only methods
Data augmentation is often employed to produce LQ images since they can provide similar domain information as the LQ evaluation datasets [11, 10, 9]. However, these on-the-fly augmentation methods usually degrade the discriminability of the images and make them hard to be identified in model training. Recent works discover the relationship between the feature norm and the image quality [19, 15]. With the awareness of image qualities, the loss functions are designed to assign different importance to the images according to their qualities. They attempt to improve the performance on the LQ images while keeping the high accuracy on the HQ images. Nevertheless, the incompatibility between the LQ data and the classification training process prohibits the encoder from effectively learning the hard LQ images. PENB [20] forces the alignment between the distribution of LQ and HQ data in a global view, which may suffer from such a problem. Instead of learning LQ data with classification-based method, we introduce contrastive learning to better understand them with the instance-level supervision. Furthermore, by quality partitioning, we can prevent the LQ images from misleading the classification-based HQ learning.
@@ II-B Quality-invariant Methods
To deal with the domain gap between the LQ and the HQ face images, some methods attempt to transfer the knowledge from the teacher model trained with the HQ images to the student model which focuses on the LQ images. In [11, 10], a pretrained teacher network is frozen and then the student network is trained to extract features similar to the teacher with the down-sampled HQ images. TCN [9] takes a contrastive loss to constrain the relationship between the HQ and the LQ images. RAN [13] proposes multi-resolution Generative Adversarial Networks (GAN) to synthesize the cross-resolution images and a feature adaption network to enhance the communication between the feature extractors for the HQ and the LQ images. DDL [14] splits the data into two subsets according to the predefined semantic difficulty including yaw, race, and distance during image capture. A similarity-based KL-divergence loss is applied on the two subsets.
These methods pursue the knowledge transfer from the HQ to the LQ images. However, with such a single-way learning process, the performances of the learned models on the HQ images are stumbled. Our QGFace partitions the training data according to their qualities and applies different supervising methods on them, which not only suits for the LQ face recognition, but also keeps the ability of recognizing the HQ images.
@@ II-C Synthesis-based Methods
To generalize the high performance of existing high-quality face recognition to different qualities, two straightforward ways are to super-resolve the LQ images for recognition [7, 8, 21, 6] and to synthesize meaningful multi-quality training data. As they can only learn the mapping between the artificially down-sampled images and the original one, the training of generation model usually needs a well-trained feature extractor. In [7, 8, 22, 21, 6, 23, 24, 25], a pretrained feature extractor is taken to supervise the identities of the synthetic images. In [26], a GAN is trained to learn how to downgrade the high-resolution images to generate realistic low-resolution images. In [27], a DCR model is taken to reserve the identity information, while in [28, 29] the network structure is modified to improve the generation quality. Owing to the absence of large-scale labeled LR face datasets, the approaches in [30, 31] synthesize low-quality images from the high-quality dataset and make it possible to train a mixed-quality FR model. [24] only focuses on a specific type of LQ images which is the partial face images. [25] reconstructs the HQ face images with singular value decomposition to improve the reliability.
These methods could obtain satisfactory HQ images to mitigate the effect of resolution difference or lead the model to learn to discriminate LQ face images. However, as an extra generative module is needed, the overall framework is not end-to-end and computationally expensive. Besides, a well-trained feature extractor is still needed in these methods, which highly dominates the training of the synthesis-based model and the performance of the recognition process. By directly training a single feature extractor which is applicable to both the HQ and the LQ data in an end-to-end way, our method is more elegant and orthogonal with the synthesis-based methods.
@@ III Method
Fig. 1(b) illustrates the framework of our quality-guided joint training for mixed-quality face recognition. We first partition the features into HQ and LQ subsets according to the image qualities, and learn them with classification loss and contrastive loss respectively. Then, we propose a proxy-updated real-time contrastive queue to support the face identification among LQ images.
@@ III-A Quality Partitioning
Considering the domain gap between the HQ and the LQ images, we partition the mixed-quality images into LQ and HQ subsets during training. In [19, 15], it is found that the feature norm is highly correlated with the image quality. In our approach, we use the feature norm as the quality indicator for data partitioning.
@@
where $c$ is a scale parameter to keep most values of $\widehat{z}$ fall in (-1,1). The gradient flow of $\widehat{z}$ is stopped during backpropagation to keep the norm from being optimized directly.
We partition the input mixed-quality data into HQ and LQ subsets based on the quality indicator $\widehat{z}$ using a quality threshold $b$. Classification loss and contrastive learning are then applied for the HQ and the LQ feature learning respectively. We define dependent but different quality partitioning strategies for these two methods according to their inherent properties. For classification loss $L_{class}$, each HQ feature is processed individually. We apply the quality comparison on the features in the same way, i.e. the features with quality indicator $\widehat{z}_{i}$ higher than $b$ are fed into the classification loss. Contrastive loss $L_{contra}$ focuses on the relationship between sample pairs. We take the representations of the augmented image and the original one as a positive contrastive pair ($q$, $k$) for learning. We define the quality indicator of the $p-{th}$ positive pair as the lower one of its two components. The pairs whose quality indicators are lower than the quality threshold $b$ are fed into the contrastive learning. Specifically,
@@
$\mathcal{L}_{classification}$ =$=\left\{\begin{array}[]{lr}{L}_{class},&\widehat{z}_{i}>b\\ 0,&else\end{array}\right.,$ (5)
$\mathcal{L}_{contrastive}$ =$=\left\{\begin{array}[]{lr}{L}_{contra},&\widehat{z}_{p}\leq b\\ 0,&else\end{array}\right.,$ (8)
@@ III-B Quality-guided joint training
The quality-invariant methods learn the LQ information from the corresponding features of the HQ images encoded by a pretrained model, and apply contrastive loss or the KL-divergence loss on all data. Different from them, we guide the model to learn the HQ and the LQ images simultaneously by applying classification loss and contrastive learning on different features according to their qualities. In this way, the supervision compromise between the HQ and the LQ features is minimized, and we can make full use of images with different qualities to train the encoder for the mixed-quality face recognition task. During the evaluation process, these two loss functions and their related structures will be dropped, and only a single feature extractor is reserved which can process images without distinction of qualities to enable mixed-quality face recognition.
@@ III-B1 Classification loss for HQ features
Margin-based softmax loss functions are widely used in face recognition and have achieved SOTA performances on the HQ evaluation datasets [2, 1, 3]. This kind of methods add a margin parameter to enlarge the distance between the matched image-identity pair and the unrelated ones to discriminate between different identities. Their loss could be shown as:
${L}=-\log\frac{e^{s\cdot f(\boldsymbol{h}_{i},\boldsymbol{W}_{y_{i}})}}{e^{s\cdot f(\boldsymbol{h}_{i},\boldsymbol{W}_{y_{i}})}+\sum^{n}_{j=1,j\neq y_{i}}e^{s\cdot f^{\prime}(\boldsymbol{h}_{i},\boldsymbol{W}_{j})}},\$ (10)
@@
With further analysis on the gradient equations of ${L}$, a gradient scaling term (GST) ${g}$ could be extracted to understand the effects of the cross-entropy softmax-based loss functions,
@@
where $P_{y_{i}}^{(i)}$ denotes the probability of an input $\boldsymbol{x}_{i}$ on its target class $y_{i}$. After substituting Equations 13 and 14 into Equation 10, we get the AdaFace loss which is taken as our classification loss $L_{class}$ for the HQ feature learning.

### v2

# QGFace: Quality-Guided Joint Training for Mixed Quality Face Recognition

@@ Abstract
The quality of a face crop in an image is decided by many factors, such as camera resolution, distance, and illumination condition. This makes the discrimination of face images with different qualities a challenging problem in realistic applications. However, most existing approaches are designed specifically for high-quality (HQ) or low-quality (LQ) images, and the performances would degrade for the mixed-quality images. Besides, many methods ask for pre-trained feature extractors or other auxiliary structures to support the training and the evaluation. In this paper, we point out that the key to better understand both the HQ and the LQ images simultaneously is to apply different learning methods according to their qualities. We propose a novel quality-guided joint training approach for mixed-quality face recognition, which could simultaneously learn images of different qualities with a single encoder. Based on quality partition, the classification-based method is employed for HQ data learning. Meanwhile, for LQ images that lack identity information, we learn them with self-supervised image-image contrastive learning. To effectively catch up the model update and improve the discriminability of contrastive learning in our joint training scenario, we further propose a proxy-updated real-time queue to compose the contrastive pairs with features from the genuine encoder. Experiments on the low-quality datasets SCface and Tinyface, the mixed-quality dataset IJB-B, and five high-quality datasets demonstrate the effectiveness of our proposed approach in recognizing face images of different qualities.
@@ I INTRODUCTION
Existing face recognition (FR) methods have achieved near perfect performance on high-quality (HQ) face images [5, 41, 24]. However, their performances significantly degrade when being applied to low-quality (LQ) images. Some works focus on face recognition, where LQ images are adopted for matching between the given probe images and the gallery images [19, 3, 20, 47, 28, 14]. However, due to the domain gap between the HQ and the LQ images, the discriminability for HQ images is somewhat reduced. It is difficult to effectively learn from both LQ and HQ images simultaneously. In reality, the quality of a face image used for recognition is influenced by numerous factors, such as camera resolution, distance, occlusion, and illumination condition. For instance, a high-resolution smart phone camera could produce low-quality face crops when there are too many faces in the picture or the photographer focuses on the scenery in the background. Meanwhile, in surveillance situations, the image qualities also vary when the distance is changing (the setting in the SCface [7] dataset). Domain gaps also exist among these low-quality images. In some limited situations where the image qualities are fixed, it is possible to select the appropriate model according to the image qualities. However, for applications such as photo albums and complicated surveillance systems with various cameras, the quality of face images vary considerably. Existing methods that focus only on HQ or LQ images are not suitable for such scenarios.
@@
The problems with mixed-quality face recognition using a single feature extractor lie in the following aspects. First, due to the domain gap between the HQ and the LQ data, it is hard to learn from and inference with images of different qualities simultaneously. Most existing methods emphasize either the HQ or the LQ images, while miss the other one. Some quality-invariant methods [47, 14, 28, 6, 11] attempt to understand the LQ images with the guidance of a pretrained HQ model. This explicitly teaches the model to lean to the LQ images and thus the performance degrades for the HQ images. Besides, as the HQ images are encoded by a fixed model, the compromise between images of difference qualities is also under-optimized. Second, we argue that there is a semantic problem when applying the classification-based method on LQ images. The identities of faces are hard to predict due to the natural variations of their poses, ages, and makeups for the HQ images. Compared with the HQ images, the LQ images typically have fewer details, which makes it challenging for the model to accurately classify them. This confusion arises because the model may struggle to differentiate between natural variations in the images and disturbances caused by the lower image qualities. As a result, LQ images usually stumble the performance of the feature extractor trained with classification loss [17].
@@
To address the above problems, we propose a novel approach, namely quality-guided joint training for mixed-quality face recognition (QGFace). Our goal is to learn both the HQ and the LQ face images simultaneously, and to recognize face images without distinction of qualities with a single feature extractor. To effectively utilize the data of different qualities, we design a quality partitioning strategy and apply different supervision signals according to the qualities of the images. Specifically, for the HQ images, the classification-based method is applied as they have achieved near-perfect performance for face recognition [41, 17, 5, 45]. For the LQ images that could mislead the classification loss, we adopt self-supervised contrastive learning, which learns the face images at the instance level and alleviates the requirements for the image details. By taking advantage of the classification methods and contrastive learning, we get an encoder which is applicable for images with different qualities. Furthermore, with the joint training strategy, we only need a single feature extractor during training and evaluation. This makes our method highly extensible, which could serve as the base component for future works.
@@
Nevertheless, taking contrastive learning from the unsupervised learning into the joint training scenario is non-trivial. Face identification asks for searching the related identity of a given image. Classification-based methods contain thousands of image-identity pairs to learn discriminative features which could meet the above requirement. For contrastive learning, it usually takes a large number of sample pairs to improve stability and accuracy. Momentum queue is a common choice [8]. It generates representations by an extra encoder which is momentum-updated with the training encoder. However, it loses track of the model which is also guided by the classification supervision signal, and the produced features are whitened as the encoder keeps an average of all the past model statuses. Inspired by BroadFace [18], we propose a novel proxy-updated real-time contrastive queue to supply positive pairs from the genuine model with a large number of negative samples which is dynamically updated. The proxy is represented by the corresponding weights of an identity in the classifier.
@@
  • We propose a quality-guided joint training strategy for face recognition (QGFace) on the mixed-quality images. To better understand the LQ images that could mislead the classification methods, we partition the images into the HQ and LQ subsets. The classification loss and contrastive loss are applied in the two subsets, respectively. In this way, we can learn comprehensively from both the LQ and the HQ faces to train a single encoder.
  • To support feature search in face identification, we propose a proxy-updated real-time contrastive queue to ensure the discriminability of the LQ images learned by contrastive learning. The identity proxy representations are taken to update the queue features encoded in the previous training step.
  • We conduct extensive experiments on high, low, and mixed quality datasets. Our approach achieves competitive results on different categories of datasets compared with the SOTA methods, which demonstrates the effectiveness of our QGFace.
Figure 1: Conventional pure classification method vs. our QGFace. (a) An FR training pipeline with a single classification loss. All input images are taken into a single main loss. (b) Our pipeline additionally takes contrastive loss for low-quality images. We apply data augmentation to images and split their features into two parts: HQ and LQ ones. The HQ images are supervised by the classification loss, whereas the LQ images and their related HQ images are sent to the contrastive function. The gradient flow of the HQ data is stopped in contrastive learning. A real-time queue is designed to provide an effective feature queue and support the large-scale feature comparison.
@@ II-A Encoder-only methods
Data augmentation is often used to produce LQ images, as they can provide domain information similar to the LQ evaluation datasets [14, 28, 47]. However, these on-the-fly augmentation methods generally degrade the discriminability of images and make it difficult to identify them in model training. Recent work discovers the relationship between the feature norm and image quality [29, 17]. With awareness of the qualities of the image, the loss functions are designed to assign different importance to the images according to their qualities. They try to improve the performance of the LQ images while maintaining high accuracy on the HQ images. However, the incompatibility between the LQ data and the classification training process prevents the encoder from effectively learning the hard LQ images. PENB [22] forces the alignment between the LQ and HQ data distribution from a global point of view, which may suffer from this problem. Instead of learning LQ data with the classification-based method, we introduce contrastive learning to better understand them with the instance-level supervision. Furthermore, by quality partitioning, we can prevent LQ images from misleading classification-based HQ learning.
@@ II-B Quality-invariant Methods
To deal with the domain gap between the LQ and the HQ face images, some methods attempt to transfer the knowledge from the teacher model trained with the HQ images to the student model which focuses on the LQ images. In [14, 28], a pretrained teacher network is frozen and then the student network is trained to extract features similar to the teacher with the down-sampled HQ images. TCN [47] takes a contrastive loss to constrain the relationship between the HQ and the LQ images. RAN [6] proposes multi-resolution Generative Adversarial Networks (GAN) to synthesize the cross-resolution images and a feature adaption network to enhance the communication between the feature extractors for the HQ and the LQ images. DDL [11] divides the data into two subsets according to the predefined semantic difficulty, including yaw, race and distance during image capture. A similarity-based KL-divergence loss is applied to the two subsets.
These methods pursue the transfer of knowledge from the HQ to the LQ images. However, with such a single-way learning process, the performances of the learned models on the HQ images are stumbled. Our QGFace partitions the training data according to their qualities and applies different supervising methods on them, which not only suits for the LQ face recognition, but also keeps the ability of recognizing the HQ images.
@@ II-C Synthesis-based Methods
To generalize the high performance of existing high-quality face recognition to different qualities, two straightforward ways are to super-resolve the LQ images for recognition [3, 20, 48, 19] and to synthesize meaningful multi-quality training data. As they can only learn the mapping between the artificially downsampled images and the original one, the training of generation model usually needs a well-trained feature extractor. In [3, 20, 44, 48, 19, 4, 34, 13], a pre-trained feature extractor is taken to supervise the identities of the synthetic images. In [1], a GAN is trained to learn how to downgrade high-resolution images to generate realistic low-resolution images. In [32], a DCR model is taken to preserve identity information, while in [39, 38] the structure of the network is modified to improve the quality of the generation. Due to the absence of large-scale labeled LR face datasets, the approaches in [23, 25] synthesize low-quality images from the high-quality dataset and make it possible to train a mixed-quality FR model. [34] only focuses on a specific type of LQ images, which are partial face images. [13] reconstructs the HQ face images with singular value decomposition to improve reliability.
These methods could obtain satisfactory HQ images to mitigate the effect of resolution difference or lead the model to learn to discriminate LQ face images. However, since an additional generative module is needed, the overall framework is not end-to-end and computationally expensive. Besides, a well-trained feature extractor is still needed in these methods, which highly dominates the training of the synthesis-based model and the performance of the recognition process. By directly training a single feature extractor which is applicable to both the HQ and the LQ data in an end-to-end way, our method is more elegant and orthogonal with the synthesis-based methods.
@@ III Method
Fig. 1(b) illustrates the framework of our quality-guided joint training for mixed-quality face recognition. We first partition the features into HQ and LQ subsets according to the image qualities, and learn them with classification loss and contrastive loss, respectively. Then, we propose a proxy-updated real-time contrastive queue to support the face identification among LQ images.
@@ III-A Quality Partitioning
Taking into account the domain gap between the HQ and the LQ images, we partition the mixed quality images into the LQ and HQ subsets during training. In [29, 17], it is found that the feature norm is highly correlated with the image quality. In our approach, we use the feature norm as the quality indicator for data partitioning.
@@
where $c$ is a scale parameter to keep most values of $\widehat{z}$ in (-1,1). The gradient flow of $\widehat{z}$ is stopped during backpropagation to keep the norm from being optimized directly.
We partition the mixed-quality input data into the HQ and LQ subsets based on the quality indicator $\widehat{z}$ using a quality threshold $b$. Classification loss and contrastive learning are then applied for the HQ and the LQ feature learning, respectively. We define dependent but different quality partitioning strategies for these two methods according to their inherent properties. For classification loss $L_{class}$, each HQ feature is processed individually. We apply the quality comparison to the features in the same way, i.e., the features with the quality indicator $\widehat{z}_{i}$ higher than $b$ are fed into the classification loss. The contrast loss $L_{contra}$ focuses on the relationship between the pair of samples. We take the representations of the augmented image and the original one as a positive contrastive pair ($q$, $k$) for learning. We define the quality indicator of the pthp-{th} positive pair as the lower one of its two components. Pairs whose quality indicators are lower than the quality threshold $b$ are fed into contrastive learning. Specifically,
@@
$\mathcal{L}_{classification}$ $=\left\{\begin{array}[]{lr}{L}_{class},&\widehat{z}_{i}>b\\ 0,&else\end{array}\right$., (5)
$\mathcal{L}_{contrastive}$ $=\left\{\begin{array}[]{lr}{L}_{contra},&\widehat{z}_{p}\leq b\\ 0,&else\end{array}\right$., (8)
@@ III-B Quality-guided joint training
The quality-invariant methods learn the LQ information from the corresponding features of the HQ images encoded by a pretrained model, and apply contrastive loss or the KL-divergence loss on all data. Differently from them, we guide the model to learn the HQ and the LQ images simultaneously by applying classification loss and contrastive learning on different features according to their qualities. In this way, the supervision compromise between the HQ and the LQ features is minimized, and we can make full use of images with different qualities to train the encoder for the mixed-quality face recognition task. During the evaluation process, these two loss functions and their related structures will be eliminated and only a single feature extractor is reserved, which can process images without distinction of qualities to enable mixed-quality face recognition.
@@ III-B1 Classification loss for HQ features
Margin-based softmax loss functions are widely used in face recognition and have achieved SOTA performances on the HQ evaluation datasets [41, 5, 24]. This kind of method adds a margin parameter to enlarge the distance between the matched image-identity pair and the unrelated ones to discriminate between different identities. Their loss could be shown as:
${L}=-\log\frac{e^{s\cdot f(\boldsymbol{h}_{i},\boldsymbol{W}_{y_{i}})}}{e^{s\cdot f(\boldsymbol{h}_{i},\boldsymbol{W}_{y_{i}})}+\sum^{n}_{j=1,j\neq y_{i}}e^{s\cdot f^{\prime}(\boldsymbol{h}_{i},\boldsymbol{W}_{j})}},\\$ (10)
@@
With further analysis on the gradient equations of L{L}, a gradient scaling term (GST) g{g} could be extracted to understand the effects of the cross-entropy softmax-based loss functions,
@@
where $P_{y_{i}}^{(i)}$ denotes the probability of an input $\boldsymbol{x}_{i}$ on its target class $y_{i}$. After substituting 13 and 14 into 10, we get the AdaFace loss, which is taken as our classification loss $L_{class}$ for the HQ feature learning.